OpenStudy (anonymous):

2x2 + x - 4 = 0 how many solutions are there

OpenStudy (mathstudent55):

Us the discriminant of the quadratic formula. b^2 - 4ac is the discriminant. In 2x^2 + x - 4 = 0, a = 2, b = 1, and c = -4 b^2 - 4ac = 1^2 - 4(2)(-4) = 1 + 32 = 33 If the discriminant is positive, there are two real roots. If the discriminant is zero, there is one real root. If the discriminant is negative, there are two negative roots.

OpenStudy (anonymous):

so there two reals roots...i got 33 to

OpenStudy (mathstudent55):

Right, since the discriminant is a positive number, there are two different real roots.